OpenStudy (anonymous):

A Web music store offers two versions of a popular song. The size of the standard version is 2.1 megabytes (MB). The size of the high-quality version is 4.2 MB. Yesterday, there were 1020 downloads of the song, for a total download size of 2982 MB. How many downloads of the high-quality version were there?

OpenStudy (one098):

Easy. They tell you the standard version is 2.1, the size of quality version is 4.2. If yesterday there were a total download size of 2982, subtract 1020 from that.

OpenStudy (anonymous):

easy for you.. :)

OpenStudy (one098):

Well... sorry about that, but did you get your answer?

OpenStudy (anonymous):

no still figuring out the equation to do it.

OpenStudy (one098):

Oh, but subtract 2982-1020 and that would be your answer.

OpenStudy (anonymous):

1962

OpenStudy (one098):

Yes.

OpenStudy (anonymous):

then how do I figure out how many were high quality?

OpenStudy (anonymous):

Just make a linear system, let x be the number of downloads of standard version and y for the high-quality version

OpenStudy (anonymous):

Let X be the number of times small file size is downloaded Let Y be the number of times big file size is downloaded 2.1X+4.2Y=2982 X+Y=1020 Do you agree with that series of equations?

OpenStudy (anonymous):

yes I got that part..

OpenStudy (anonymous):

now just solve the linear system AMelch8 showed

OpenStudy (anonymous):

400

OpenStudy (anonymous):

So you are saying Y=400? Let's plug them into the equations and check: If Y= 400, from the second equation X+Y=1020 therefore X= 620. Now to check the first equation: 2.1(620)+4.2(400)=2982, this satisfies your two equations. Good work. It's always a good idea to show your work and then check it after. The good thing about linear algebra is it always gives you a chance to independently check your answer.